A 1.6 g samlpe of metal that appears to be gold required 5.8 j of energy to make an 18 degrees celcius temerature change. is the metal pure gold?
1
Expert's answer
2012-06-08T07:58:52-0400
Q = m*s*dT
m(metal) = 1.6 g
dT = 18 C
s(metal) = Q/(m*dT) = 5.8 J /
(1.6 g * 18 C) = 0.201 J/(g*C)
The specific heat of gold is 0.13 J/(g*C).

Therefore the metal cannot be pure gold.